Passengers in a whirly-go-round (a "puke your guts out" amusement ride) are subjected to accelerations of 2.5 g's at the end of a 3.5 m pole as they move in a circular path. What speed (in m/s) do they experience?
Answer #1

a = v^2/r

2.5* 9.8 = v^2/ 3.5

v = 9.26 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck